OpenStudy (anonymous):

Find the point on the terminal side of θ = -3pi/4 that has an x coordinate of -1.

OpenStudy (anonymous):

terminal side of what ??

OpenStudy (anonymous):

I assume the equation itself.

OpenStudy (tkhunny):

Okay, so \((-1,Something)\) is at \(-3\pi/4\). 1) Where is \(-3\pi/4\)? What quadrant is it in? 2) What is the reference angle there? 3) You've a Pythagorean theorem to help you the rest of the way.

OpenStudy (anonymous):

It's in the second quadrant. Reference angle?

OpenStudy (tkhunny):

Are you sure you have the background to solve this problem? It's in the 3rd quadrant. \(+3\pi/4\) is in the 2nd quadrant. It is not encouraging that you don't know what a reference angle is. You will need that information. Draw a ray from the origin in the direction of \(\3\pi/4\). What angle does this ray make with the x-axis?

OpenStudy (anonymous):

Gotcha because it is negative. Well below the 3pi/4 is 225 degrees? This is from the unit circle.

OpenStudy (tkhunny):

@tanjeetsarkar96 Are you just guessing? You had it right before, but please do not just post answers.

OpenStudy (tkhunny):

Kind of. It's really -135º, but this gest you to the same place as +225º. Good work. Now the reference angle.

OpenStudy (anonymous):

Well the answer does make since now that I think about it. Since the unit circle is only 1 digit for the x axis and the y axis, going to the very edge would force the y corrdinate to be 0. Is this the correct reasoning?

OpenStudy (anonymous):